OpenStudy (nottim):

Water is at temp. 23 deg C and air temp in car is -7 degrees C. This can be modelled by T(t)=-7+Ae^(kt), degrees C, t hrs after the car door closes. What is A?

OpenStudy (perl):

T(0) = 23, so plug in t = 0 and solve for A

OpenStudy (perl):

-7 + Ae^(k*0) = 23 -7 + A*1 = 23 A = 23 + 7 A = 30

OpenStudy (nottim):

Oh whoops. Thanks.

OpenStudy (perl):

and you're not Tim? what should i call you

OpenStudy (nottim):

The Great Not. Or NotTim the Spectacular.

OpenStudy (perl):

ok :)